OpenStudy (peterg2001):

help please will medal Choose the answer based on the most efficient method as presented in the lesson. If the first step in the solution of the equation 2x - 8 = 5x + 3 is "subtract 2x," then what should the next step be? subtract 5x add 8 subtract 3

OpenStudy (peterg2001):

@haleyelizabeth2017 @AlexandervonHumboldt2

pooja195 (pooja195):

What do you think?

pooja195 (pooja195):

After subtracting 2x you are left with 2x - 8 = 5x + 3 -8=3x+3

OpenStudy (haleyelizabeth2017):

2x - 8 = 5x + 3 If you subtract 2x, you'd be left with -8=3x+3 you always want something on both sides of the equation, so what would make sense? add 8 or subtract 3?

OpenStudy (haleyelizabeth2017):

Let me show you this... adding eight leaves you with: 0=3x+11 Subtracting 3 leaves you with: -11=3x. Think you can figure out which one is the most logical answer?